20170523 » History » Version 1
Boyd, Steven, 22 May 2017 09:51
1 | 1 | Boyd, Steven | h1. Ops/Comm Scheduling meeting : 23/5/2017 |
---|---|---|---|
2 | |||
3 | Location : CR21 at 14:00 pm |
||
4 | |||
5 | Phone access : *New* phonebridge number |
||
6 | UK: 08082380274 |
||
7 | US: 631.621.5253 |
||
8 | Access code: 1779339517 |
||
9 | |||
10 | h2. Agenda |
||
11 | |||
12 | * Actions |
||
13 | ** -Macwaters - Re-make dodgy compressor2 remote cable connector - stands- |
||
14 | ** Macwaters - Remove old overheated CSW71 to R9 _waiting for John Govans and the van_ - in progress |
||
15 | ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress |
||
16 | ** Macwaters - Report on status of FC psu repair (stands) |
||
17 | *** The spare is installed and has been run into a short |
||
18 | ** Hall Manager - water system change-over procedure (stands) - _collar people on Friday_ |
||
19 | ** EOverton - Status of @miceraid5@ disk cleanup (stands) _now has a script but needs to run it_ |
||
20 | ** Progress on Hall list [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) _DL were down on masse and made significant progress. Need to confirm if any remaining items_ |
||
21 | ** CWhyte to request survey of TOF1 - _requested, next week as Tony off this week_ |
||
22 | ** SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings |
||
23 | ** AKurup to check which settings are being used by the active Alarm Handlers |
||
24 | ** SBoyd to conscript extra shift cover for the end of next week : Done. Need BLOCs for the 31/1 and 2 of June (i.e. end of cycle) |
||
25 | |||
26 | * DC update |
||
27 | * MOM update |
||
28 | * The tracker saga |
||
29 | * Hydrogen system update |
||
30 | * Run plan |
||
31 | ** Tracker may be ready to operate on Friday |
||
32 | *** Is there still Helium in the tracker? |
||
33 | ** LH2 running will not happen this cycle |
||
34 | ** Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17); |
||
35 | ** Magnet commissioning: Sunday (28May17)—Monday (29May17); and |
||
36 | ** Empty-channel data taking (Monday 29May17)—Friday (02Jun17). |
||
37 | * AOB |
||
38 | |||
39 | * Last meeting : [[20170516|Tuesday May 9 2017]] |
||
40 | * Next meeting : [[20170530|Tuesday May 30 2017]] |
||
41 | |||
42 | h3. Minutes |
||
43 | |||
44 | * Actions |
||
45 | ** -Macwaters - Re-make dodgy compressor2 remote cable connector - stands- |
||
46 | ** Macwaters - Remove old overheated CSW71 to R9 _waiting for John Govans and the van_ - in progress |
||
47 | ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress |
||
48 | ** Macwaters - Report on status of FC psu repair (stands) |
||
49 | *** The spare is installed and has been run into a short |
||
50 | ** Hall Manager - water system change-over procedure (stands) - _collar people on Friday_ |
||
51 | ** EOverton - Status of @miceraid5@ disk cleanup (stands) _now has a script but needs to run it_ |
||
52 | ** Progress on Hall list [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) _DL were down on masse and made significant progress. Need to confirm if any remaining items_ |
||
53 | ** CWhyte to request survey of TOF1 - _requested, next week as Tony off this week_ |
||
54 | |||
55 | * New actions |
||
56 | ** SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings |
||
57 | ** AKurup to check which settings are being used by the active Alarm Handlers |
||
58 | ** SBoyd to conscript extra shift cover for the end of next week |